Tohban Report 2017-04-19
Tohban Reports | |
---|---|
Start Date: | 29 March 2017 |
End Date: | 5 April 2017 |
Tohban: | Milo Buitrago-Casas |
Tohban email: | milo@ssl.berkeley.edu |
Next Tohban: | Jim McTiernan |
List all reports |
Solar Activity
The Sun was active this week, producing a handful of B-class flares and two C-flares, one yesterday and one this morning. Currently, there is one AR peeping out on the Est limb and is expected to remain on disk this coming week with a slight chance to host M-class flares.
How many GOES flares occurred?
Flares above B, C, M, X class were 18 1 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 8 1 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 0 0 0 0
There were RHESSI flares/GOES flares 22 / 19 over the time range 11-Apr-17 18-Apr-17
Memory Management
Early this week, the SSR was quite high ~40% at due to the four detectors that were on last week. After turning off detectors 1 and 8 on April 12, the SSR was gradually emptying reaching a minimum of 9% at the end of the day on Monda April 17. In response to the rotation on the disk of a big Active Region, detectors 1 and 8 were turned on again so the SSR is expected to gradually increase this week. Currently, it is at 18.4%.
Spacecraft Status
Detectors 3 and 6 are recording events. Due to a big active region on the solar disk, detectors 1 and 8 were turned on.
The cold tips are 127.3 and 125.1 K as of 19-Apr-2017. The cold plates, 1 and 2 are 152.5 and 151.0 K. Cryocooler power is steady at 74.9W. See:
http://sprg.ssl.berkeley.edu/~ayshih/soh/latest_temperatures.pdf
for a summary plot of recent temperatures, efficiencies and accelerometer values.
Data Gaps
Relatively large (>300 sec) data gaps as follows:
GAP START TIME GAP END TIME GAP (SEC) 2017-04-15T17:25:00.000 -- 2017-04-15T22:40:00.000 18900.000 2017-04-15T17:25:00.000 -- 2017-04-15T22:35:00.000 18600.000 2017-04-13T13:50:00.000 -- 2017-04-13T14:00:00.000 600.00000
Detector issues
Because the moderate solar activity early this week, and the presence of a new active region on the Est solar limb, it was decided to turn on detectors 1 and 8 the past Monday Abril 17 around 21:30UT. Temperatures on the cold plates and tips looked stable during the last couple of days after turning on those two detectors. Also the SSR filling level has been rather low (~15% at the end of the passes).
Commanding were the follow:
2017-107-21:31:18 start idib_on_g1 2017-107-21:31:35 start idib_on_g8 2017-107-21:32:05 start ihv_ramp_stage2_g1 ; 1000 V 2017-107-21:32:20 start ihv_ramp_stage2_g8 ; 1000 V
17-107-23:09:42 start ihv_ramp_stage3_g1 17-107-23:09:58 start ihv_ramp_stage3_g8
17-107-23:14:37 start sc_ssr_rtime halted
17-108-00:40:32 start idib_chg_thrshld(8, FRONT, FAST, 0x90) 17-108-00:40:36 /IDPUDUMPTABL TABLE=DIBTBL8 17-108-00:40:36 /IDPUTABLE8 OFFSET=FRONTFASTDAC 17-108-00:41:06 /IDPULOAD VALUE=0x90
17-108-00:43:46 start idib_chg_thrshld(1, FRONT, FAST, 0x80) 17-108-00:43:49 /IDPUDUMPTABL TABLE=DIBTBL1 17-108-00:43:49 /IDPUTABLE1 OFFSET=FRONTFASTDAC 17-108-00:43:58 /IDPULOAD VALUE=0x80
17-108-00:45:22 start idib_chg_thrshld(1, REAR, FAST, 0x60) 17-108-00:45:26 /IDPUDUMPTABL TABLE=DIBTBL1 17-108-00:45:27 /IDPUTABLE1 OFFSET=REARFASTDAC 17-108-00:45:34 /IDPULOAD VALUE=0x60
17-108-00:49:52 start sc_ssr_rtime halted
Detectors 1, 3, 6 and 8 are on and working. There is a new active region on disk (AR2651) that produced a C-class flare this morning and will remain active for the next couple of days. So it was determined to keep the detectors 1 and 8 on unless there is an issue with the temperatures.
Notes
There was an explained attenuator change on 13 April 2017 at 13:55 UT. Albert's best guess is that this is (at least partially) corrupted data. At first glance, there is not any other explanation for why the front-decimation level could vary so wildly on short timescales, from minimal decimation to throwing out ~255/256 of counts. In the corresponding lightcurve plot one orbit earlier, the decimation level also appears to change during eclipse, which is weird as well. Further investigation is needed.
Spacecraft Management
Decimation | Active/Vigorous |
HLAT Decimation | Rear decimation |
Night time data (fronts) | plus/minus 4 minutes from daylight |
Night time data (rears) | plus/minus 4 minutes from daylight |
Require extra passes? | No |
Requirement for moving pointer? | No |
Attenuator operation | None |
Detector problems? | See notes above. |